# Investigating Adversarial Vulnerability and Implicit Bias through Frequency Analysis

### v2

@@ Abstract
footnotetext:
Correspondence to: fabio.anselmi@units.it and alejandro.rodriguezgarcia@units.it.
Code is available at https://github.com/lorenzobasile/ImplicitBiasAdversarial
@@
Despite their impressive performance in classification tasks, neural networks are known to be vulnerable to adversarial attacks, subtle perturbations of the input data designed to deceive the model. In this work, we investigate the relation between these perturbations and the implicit bias of neural networks trained with gradient-based algorithms. To this end, we analyse the network’s implicit bias through the lens of the Fourier transform. Specifically, we identify the minimal and most critical frequencies necessary for accurate classification or misclassification respectively for each input image and its adversarially perturbed version, and uncover the correlation among those. To this end, among other methods, we use a newly introduced technique capable of detecting non-linear correlations between high-dimensional datasets. Our results provide empirical evidence that the network bias in Fourier space and the target frequencies of adversarial attacks are highly correlated and suggest new potential strategies for adversarial defence.
1 Introduction
@@
Regardless of their overparameterisation, deep neural networks can avoid over-fitting and generalise well to new data. This surprising property has been attributed to the so-called implicit bias of the network, which refers to the ability of the gradient to converge to solutions with good generalisation properties among the many that correctly label the training data [1, 2, 3]. Understanding this bias has been the focus of extensive research in recent years. However, it remains unclear how it leads to generalisation in non-linear complex neural networks. Indeed, with the exception of simple models [4], a formal characterisation of the implicit bias of a neural network remains a formidable challenge.
@@
Besides generalisation, another important (and related) property of neural networks is robustness, i.e. their ability to maintain performance and accuracy when exposed to noise in input data and adversarial attacks. In particular, adversarial attacks have gained significant attention since the work of [5], as they consist of perturbations of the input that are imperceptible to humans but can drastically alter the response of a trained network. Although substantial research has been conducted to understand this vulnerability [6, 7, 8, 9, 10] and numerous strategies have been developed to mitigate its effects [11, 12, 13, 14, 15], a definitive explanation of its origins remains elusive.
In this work, we explore the relationship between two crucial properties of a trained network: its vulnerability to adversarial attacks and its implicit bias. While this problem has been the focus of recent studies, explicit results have so far only been achieved for simple networks [16, 17, 18]. In this context, the research presented in [19] is particularly valuable since it offers an algorithm to investigate a specific aspect of implicit bias, even in complex networks. Their approach highlights the essential input frequencies required to maintain the accuracy of a trained network. In particular, these frequencies are computed by training a learnable modulatory mask for each input image, which filters the image’s frequency content and reduce it to the bare minimum needed to maintain the correct classification. These masks act as a fingerprint for the network’s implicit bias in Fourier space since they highlight the frequencies the network depends on when processing inputs. In this work, we adapt the methodology from [19] to investigate, for each input image, the correlation between the frequencies essential for the correct classification of the image and those targeted by adversarial attacks for its misclassification. Fig. 1 displays examples of clean and attacked images before (A, B) and after (C, D) being filtered by, respectively, essential frequency masks and adversarial frequency masks. Our primary objective is to provide empirical proof that the network bias is highly correlated with the adversarial attacks in Fourier space. However, defining and computing this correlation is challenging due to the high-dimensional nature of the modulatory mask sets and their potential highly non-linear correlation. To address these challenges, we use different correlation estimators among which a recently introduced non-linear one [20] relying on the observation that the intrinsic dimensionality ($I_{d}$) of a dataset is closely linked to the correlations among the various features defining the data points. The idea is that these correlations determine the allowed regions where the data points live, effectively shaping the underlying manifold and its intrinsic dimensionality. Our findings indicate a strong correlation between the manifolds of the feature spaces defined by the two types of masks, providing empirical evidence of the connection between the network bias in Fourier space and the target frequencies of adversarial attacks.
The main contributions of this work can be summarised as follows:
  • We provide the first detailed quantitative computation of the essential frequencies in adversarially attacked images.
  • We demonstrate a strong correlation between implicit bias and adversarial perturbations between the essential frequencies, identified as in the first contribution, of clean and adversarial images.
  • We show that the essential frequencies required to classify images in a certain class are sufficient to counteract adversarial attacks, suggesting a promising approach for designing new defense protocols.
@@
Figure 1: Examples of CIFAR-10 [21] images before and after being filtered by the Fourier masks: (A): original input images (B): adversarial images generated with $\ell_{\infty}$ Fast Minimum Norm [22] attack on ResNet-20 [23] (C): images filtered by essential frequency masks (D): adversarial images filtered by adversarial frequency masks.
2 Related work and background
2.1 Implicit Bias, Implicit Fourier Bias and Robustness
@@
The implicit bias of state-of-the-art models has been shown to play a critical role in the generalisation properties of deep neural networks [1, 3]. Research over the years has thoroughly investigated the implicit bias present in training algorithms for different network architectures. Although an explicit characterisation of the bias has been found only in simple networks such as linear[29, 30, 27, 31, 32], and some nonlinear models [33, 34, 35, 17, 36, 37, 38], the phenomenon is still lacking a final explanation. Interestingly, a study by Frei et al. [17] demonstrates that the implicit bias in neural networks acts as a "Double-Edged Sword": while it enhances generalisation properties, it can also compromise the network’s robustness, presenting both benefits and potential drawbacks. Similar conclusions are reached in [39]. Conversely, [40] reveals a deep connection between robust classification and implicit bias, suggesting that the implicit bias can improve robustness in linear regression. Finally, in [41] the authors show that a shallow network with a polynomial ReLU activation trained by gradient flow not only generalises well but is also robust to adversarial attacks. These studies highlight that, while some insight into the relationship between implicit bias and the robustness of a neural network has been gained, a clear understanding is still lacking. One interesting effect of the implicit bias in neural networks is their tendency to learn specific frequencies in the target function during training, a phenomenon known as spectral bias [42]. This bias leads the network to learn low-complexity functions, potentially explaining its ability to generalise [43, 44, 45, 46]. Additionally, the implicit bias can also significantly influence the types of input features extracted by a trained neural network. For instance, [19] demonstrates that very few image frequencies in the Fourier domain are enough for the network to perform classification. These findings have helped characterise the spectral bias of neural networks with a focus on the input space rather than the target function (as in [42]). Moreover, empirical evidence shows a strong relationship between network robustness and the statistics of the Fourier spectra of the input data [28] or architecture [47]. Detection strategies in the Fourier domain have also been used to defend against adversarial attacks [48].
2.2 Adversarial attacks
@@
Artificial Neural Networks are well known to be vulnerable to adversarial attacks [5, 6]. These attacks involve manipulating an input data point in a way that deceives an otherwise well-performing classifier, by making small alterations to a correctly classified data point. Numerous techniques have been proposed to create such adversarial examples, beginning with the Fast Gradient Sign Method (FGSM) [6], followed shortly by variants such as Projected Gradient Descent (PGD) [13]. Both these methods employ gradient information to generate an appropriate adversarial example while ensuring that the $\ell_{p}$ norm of the perturbation remains below a fixed threshold $\epsilon$. These algorithms were primarily developed for effectiveness rather than optimality, which may limit their ability to generate input samples with minimal perturbations, resulting in them being classified as "maximum confidence" attacks. In contrast, "minimum norm" attacks prioritise the identification of adversarial examples with the least amount of perturbation by minimising its norm. In this regard, some of the most notable proposals are L-BFGS [5], the Carlini and Wagner attack [49], DeepFool [50] and the more recent Fast Minimum Norm (FMN) attack [22], which seeks to combine the efficiency of FGSM and PGD with optimality in terms of perturbation norm.
@@
Figure 2: Schematic representation of the method employed to obtain essential frequency masks and adversarial frequency masks. Only one channel is displayed for visualization purposes. Full details are provided in Sec. 4.4.
3 Methods
3.1 Modulatory Masks
Our main tools for understanding the implicit spectral bias and the geometry of adversarial examples are modulatory masks. These masks are computed to extract information about the essential input frequencies that the network requires to perform a specific classification task. To generate these masks, we employ an algorithm similar to the one described in [19], as illustrated in Fig. 2. We train masks to modulate the frequency content of an image by performing element-wise multiplication between each entry of the Fast Fourier Transform (FFT) of the image and the corresponding entry of the mask. Each entry of the mask is a learnable scalar between 00 and $1$. Specifically, starting from an image $x$, we compute its FFT $\mathcal{F}x$ and multiply it element-wise with a learnable mask $M$. The mask has the same shape of the image $x$ (and its FFT $\mathcal{F}x$), meaning that if the image has RGB encoding we train a separate mask for each channel, and its entries are constrained to be in $[0,1]$. The result of this multiplication is then projected back in pixel space by taking the real part of its inverse Fourier transform, thereby obtaining a new filtered image $x_{F}$:
@@ 3.1 Modulatory Masks
The image $x_{F}$ is then fed into the trained classification model to obtain a prediction. We produce two sets of masks. The masks belonging to the first set encode the essential frequencies of an image to be correctly classified by the neural classifier, thus we will refer to these as essential frequency masks ($M_{EF}$). The second set is composed of masks that encode the essential frequency content required to maintain the effectiveness of an adversarial attack, that is, the essential frequencies needed to misclassify an adversarially perturbed image. We will refer to these masks as adversarial frequency masks ($M_{AF}$). Some examples of adversarial frequency masks are shown in Fig. 3 (the corresponding $M_{EF}$ masks are shown in the Appendix in Fig. 5). Both sets of masks are learned using a preprocessing layer attached to a trained ANN with freezed parameters. The essential frequency masks are trained by optimising the Cross-Entropy loss of the entire model (consisting of the preprocessing layer and the trained classifier) on the original samples. Conversely, for adversarial frequency masks, the training objective is the Cross-Entropy with respect to the adversarial class (to preserve misclassification), and the masks are trained on adversarial data.
The key property of the learned masks is their sparsity, which is achieved by enforcing an $\ell_{1}$ norm regularisation on the entries of the mask during training. This regularisation ensures that the mask accurately captures only the essential frequency content needed to accomplish a specific task, such as correctly classifying an input or misclassifying an adversarial example. Our primary objective is to determine whether a correlation exists between these distinct sets of masks.
Figure 3: Examples of adversarial frequency masks, represented as RGB images. The labels refer to the classification of the clean image. The masks were obtained using CIFAR-10 and the Fast Minimum Norm attack on ResNet-20.
3.2 High-dimensional correlation
@@
As mentioned earlier, to examine the statistical relationship between implicit bias and adversarial attacks, we compute the correlations between two feature spaces: that of the essential frequencies for image classification and that of the essential frequencies required for the adversarial attack to be successful. In this section we introduce the main tools that we employ to this aim. One of the most common approach for investigating correlations is based on the Pearson correlation coefficient ($R^{2}$) between variables. However, this method can only be applied to univariate datasets, making it impractical when dealing with multidimensional data, like in our case. The direct analogous to $R^{2}$ in a multivariate scenario is Canonical Correlation Analysis (CCA) [53].
@@
CCA works by linearly projecting the two multidimensional datasets to a subspace where the correlation between the projected variables is maximised. A more modern derivative of CCA is Singular Value CCA (SVCCA) [54], which employs Principal Component Analysis before CCA, with the objective of denoising and reducing the dimensionality of data. SVCCA proved to be particularly useful in analysing the internal representations of neural networks [54]. In this direction, other correlation metrics have been proposed, including Projection Weighted CCA [55], Centered Kernel Alignment [56] and the recent Intrinsic Dimension Correlation ($I_{d}$or) [20].
@@
$I_{d}$or is a method for quantifying correlation that is well suited for non-linear multidimensional spaces. It is based on the concept of Intrinsic Dimension ($I_{d}$) of data, which refers to the minimum number of variables required to describe a dataset. The key idea behind $I_{d}$or is that if there is a correlation between two datasets (i.e., the information from one can partially reconstruct the other), the combined $I_{d}$ of the datasets will be lower than the sum of their individual $I_{d}$s. This method provides both a correlation coefficient and a P-value, obtained through a permutation test, indicating the probability of no correlation.
@@
Finally, Cosine Similarity is another method for measuring similarity between high-dimensional vectors, defined as the cosine of the angle between two vectors. Unlike other methods, cosine similarity is not strictly a measure of correlation. It offers a localized measure of similarity for individual pairs of vectors rather than evaluating entire datasets (i.e., multiple samples of a random variable). Additionally, cosine similarity requires that both vectors exist in the same vector space, making it unsuitable for vectors with different dimensionalities.
4 Experimental results
4.1 Data

### v3

# Frequency maps reveal the correlation between Adversarial Attacks and Implicit Bias

@@ Abstract
Despite their impressive performance in classification tasks, neural networks are known to be vulnerable to adversarial attacks, subtle perturbations of the input data designed to deceive the model. In this work, we investigate the correlation between these perturbations and the implicit bias of neural networks trained with gradient-based algorithms. To this end, we analyse a representation of the network’s implicit bias through the lens of the Fourier transform. Specifically, we identify unique fingerprints of implicit bias and adversarial attacks by calculating the minimal, essential frequencies needed for accurate classification of each image, as well as the frequencies that drive misclassification in its adversarially perturbed counterpart. This approach enables us to uncover and analyse the correlation between these essential frequencies, providing a precise map of how the network’s biases align or contrast with the frequency components exploited by adversarial attacks. To this end, among other methods, we use a newly introduced technique capable of detecting nonlinear correlations between high-dimensional datasets. Our results provide empirical evidence that the network bias in Fourier space and the target frequencies of adversarial attacks are highly correlated and suggest new potential strategies for adversarial defence. Code is available at https://github.com/lorenzobasile/ImplicitBiasAdversarial
Index Terms:
Neural Networks, Implicit Bias, Adversarial Attacks, Fourier Transform
I Introduction
@@
Regardless of their overparameterisation, deep neural networks can avoid over-fitting and generalise well to new data. This surprising property has been attributed to the so-called implicit bias of the network, which refers to the ability of gradient descent to converge to solutions with good generalisation properties among the many that correctly label the training data [1, 2, 3]. Understanding this bias has been the focus of extensive research in recent years. However, it remains unclear how it leads to generalisation in nonlinear, complex neural networks. Indeed, with the exception of simple models [4], a formal characterisation of the implicit bias of a neural network remains a formidable challenge.
@@
Besides generalisation, another important (and related) property of neural networks is robustness, i.e., their ability to maintain performance and accuracy when exposed to noise in input data and adversarial attacks. In particular, adversarial attacks have gained significant attention since the work of [5], as they consist of perturbations of the input that are imperceptible to humans but can drastically alter the response of a trained network. Although substantial research has been conducted to understand this vulnerability [6, 7, 8, 9] and numerous strategies have been developed to mitigate its effects [10, 11, 12, 13], a definitive explanation of its origins remains elusive.
In this work, we explore the relationship between two crucial properties of a trained network: its vulnerability to adversarial attacks and its implicit bias. While this problem has been the focus of recent studies, explicit results have so far only been achieved for simple networks [14, 15, 16]. In this context, recent research [17] has developed algorithms to investigate specific aspects of implicit bias within complex networks, particularly by analysing the frequencies that are essential for accurate classifications. This approach involves training a modulatory mask for each input image, which selectively filters frequency content and reduces it to the bare minimum needed to maintain the correct classification. In doing so, trained masks encode a map of the essential frequencies required by the network. Hence, in the rest of this manuscript we will refer to them as frequency maps. In our work, we employ such frequency maps to explore, for each input, the relationship between the critical frequencies needed for accurate classification and those targeted by adversarial attacks for misclassification. These maps serve as a dual representation: they reveal the network’s implicit bias in Fourier space by highlighting the frequencies it relies on to process inputs, while also illustrating the specific frequency components used by adversarial attacks.
Fig. 1 displays examples of clean and attacked images before (A, B) and after (C, D) being filtered by, respectively, essential frequency maps and adversarial frequency maps. Our primary objective is to provide empirical proof that the network bias is correlated with the adversarial attacks in Fourier space. However, defining and computing this correlation is challenging due to the high-dimensional nature of the modulatory map sets and their potential highly nonlinear correlation. To address these challenges we use various correlation estimators, among which a recently introduced nonlinear one [18], that relies on the observation that the intrinsic dimensionality ($I_{d}$) of a dataset is closely linked to the correlations among the various features defining the data points.
@@
Figure 1: Examples of CIFAR-10 [19] images before and after being filtered by the Fourier maps: (A): original input images (B): adversarial images generated with $\ell_{\infty}$ Fast Minimum Norm [20] attack on ResNet-20 [21] (C): images filtered by essential frequency maps (D): adversarial images filtered by adversarial frequency maps.
Our findings indicate a strong correlation between the manifolds of the feature spaces defined by the two types of maps, providing empirical evidence of the similarity between the network bias in Fourier space and the target frequencies of adversarial attacks. The main contributions of this work can be summarised as follows:
  • We provide the first detailed quantitative computation of adversarial essential frequencies, which constitute a representation of the features targeted by adversarial attacks.
  • We demonstrate a strong correlation between the network’s implicit bias and the adversarial perturbations through the essential frequencies of clean and adversarial images.
  • We show that the essential frequencies required to classify images in a certain class are sufficient to counteract adversarial attacks, suggesting a promising approach for designing new defence protocols.
II Related work and background
II-A Implicit Bias, Implicit Fourier Bias and Robustness
Figure 2: Schematic representation of the method employed to obtain essential frequency maps and adversarial frequency maps. Only one channel is displayed for visualization purposes. Full details are provided in Sec. IV-D.
@@ II-A Implicit Bias, Implicit Fourier Bias and Robustness
The implicit bias of state-of-the-art models has been shown to play a critical role in the generalisation properties of deep neural networks [1, 3]. Research over the years has thoroughly investigated the implicit bias present in training algorithms for different network architectures. Although an explicit characterisation of the bias has been found only in simple networks such as linear [27, 28, 25, 29], and some nonlinear models [30, 31, 15, 32, 33], the phenomenon is still lacking a final explanation. Interestingly, a study by Frei et al. [15] demonstrates that the implicit bias in neural networks acts as a ”Double-Edged Sword”: while it enhances generalisation properties, it can also compromise the network’s robustness, presenting both benefits and potential drawbacks. Similar conclusions are reached in [34]. Conversely, [35] reveals a deep connection between robust classification and implicit bias, suggesting that the implicit bias can improve robustness in linear regression. Finally, [36] show that a shallow network with a polynomial ReLU activation trained by gradient flow not only generalises well but is also robust to adversarial attacks. These studies highlight that, while some insight into the relationship between implicit bias and the robustness of a neural network has been gained, a clear understanding is still lacking. One interesting effect of the implicit bias in neural networks is their tendency to learn frequencies from low to high in the target function during training, a phenomenon known as spectral bias [37]. This bias leads the network to learn low-complexity functions, potentially explaining its ability to generalise [38, 39, 40, 41]. Additionally, the implicit bias can also significantly influence the types of input features extracted by a trained neural network. For instance, [17] demonstrate that very few image frequencies in the Fourier domain are enough for the network to perform classification. These findings have helped characterise the spectral bias of neural networks with a focus on the input space rather than the target function, as in [37]. Moreover, empirical evidence shows a strong relationship between network robustness and the statistics of the Fourier spectra of the input data [26] or architecture [42]. In particular, certain frequency-based patterns in adversarial robustness have been identified, suggesting that low-frequency perturbations can be especially effective [43]. Furthermore, studies indicate that frequency-driven perturbations can be leveraged for adversarial attacks while remaining imperceptible to human observers [44]. Detection strategies in the Fourier domain have also been used to defend against adversarial attacks [45].
II-B Adversarial Attacks
@@
Artificial Neural Networks are notoriously vulnerable to adversarial attacks [5, 6]. These attacks involve manipulating an input data point in a way that deceives an otherwise well-performing classifier, by making small alterations to a correctly classified data point. Numerous techniques have been proposed to create such adversarial examples, beginning with the Fast Gradient Sign Method (FGSM) [6], followed shortly by variants such as Projected Gradient Descent (PGD) [12]. Both these methods employ gradient information to generate an appropriate adversarial example while ensuring that the $\ell_{p}$ norm of the perturbation remains below a fixed threshold $\epsilon$. These algorithms were primarily developed for effectiveness rather than optimality, which may limit their ability to generate input samples with minimal perturbations, resulting in them being classified as ”maximum confidence” attacks. In contrast, ”minimum norm” attacks prioritise the identification of adversarial examples with the least amount of perturbation by minimising its norm. In this regard, some of the most notable proposals are L-BFGS [5], the Carlini and Wagner attack [46], DeepFool [47] and the more recent Fast Minimum Norm (FMN) attack [20], which seeks to combine the efficiency of FGSM and PGD with optimality in terms of perturbation norm.
@@
III Methods
III-A Modulatory Maps
Our main tools for understanding the implicit spectral bias and the geometry of adversarial examples are modulatory maps. These maps are computed to extract information about the essential input frequencies that the network requires to perform a specific classification task. To generate the maps, we employ a procedure similar to the one described in [17], as illustrated in Fig. 2. We train maps to modulate the frequency content of an image by performing element-wise multiplication between each entry of the Fast Fourier Transform (FFT) of the image and the corresponding entry of the map. Each entry of the map is a learnable scalar between 00 and $1$. Specifically, starting from an image $x$, we compute its FFT $\mathcal{F}x$ and multiply it element-wise with a learnable map $M$. The map has the same shape of the image $x$ (and its FFT $\mathcal{F}x$), meaning that if the image has RGB encoding we train a separate map for each channel, and its entries are constrained to be in $[0,1]$. The result of this multiplication is then projected back in pixel space by taking the real part of its inverse Fourier transform, thereby obtaining a new filtered image $x_{F}$:
@@ III-A Modulatory Maps
The image $x_{F}$ is then fed into the trained classification model to obtain a prediction. We produce two sets of maps. The maps belonging to the first set encode the essential frequencies of an image to be correctly classified by the neural classifier, thus we will refer to these as essential frequency maps ($M_{EF}$). The second set is composed of maps that encode the essential frequency content required to maintain the effectiveness of an adversarial attack, that is, the essential frequencies needed to misclassify an adversarially perturbed image. We will refer to these maps as adversarial frequency maps ($M_{AF}$). Some examples of essential and adversarial frequency maps are shown in Fig. 3. Both sets of maps are learned using a preprocessing layer attached to a trained classifier with frozen parameters. The essential frequency maps are trained by optimising the Cross-Entropy loss of the entire model (consisting of the preprocessing layer and the trained classifier) on the original samples. Conversely, for adversarial frequency maps, the training objective is the Cross-Entropy with respect to the adversarial class (to preserve misclassification), and the maps are trained on adversarial data. The key property of the learned maps is their sparsity, which is achieved by enforcing an $\ell_{1}$ norm regularisation on the entries of the map during training. This regularisation ensures that the map accurately captures only the essential frequency content needed to accomplish a specific task, such as correctly classifying an input or misclassifying an adversarial example. Our primary objective is to generate these maps and to determine whether a correlation exists between these distinct sets of maps.
III-B High-dimensional correlation
@@
As mentioned earlier, to examine the statistical relationship between implicit bias and adversarial attacks, we compute the correlations between two feature spaces: that of the essential frequencies for image classification and that of the essential frequencies required for the adversarial attack to be successful. In this section we introduce the main tools that we employ to this aim.
@@
The most common approach for investigating correlations is based on the Pearson correlation coefficient ($R^{2}$) between variables. However, this method can only be applied to univariate datasets, making it impractical when dealing with multidimensional data, like in our case. The direct analogous to $R^{2}$ in a multivariate scenario is Canonical Correlation Analysis (CCA) [50]. CCA works by linearly projecting the two multidimensional datasets to a subspace where the correlation between the projected variables is maximised. A more modern derivative of CCA is Singular Value CCA (SVCCA) [51], which employs Principal Component Analysis before CCA, with the objective of denoising and reducing the dimensionality of data. SVCCA proved to be particularly useful in analysing the internal representations of neural networks [51].
@@
In this direction, other correlation metrics have been proposed, including Projection Weighted CCA [52], Centered Kernel Alignment [53] and the recent Intrinsic Dimension Correlation ($I_{d}$or) [18]. $I_{d}$or is a method for quantifying correlation that is well suited for nonlinear multidimensional spaces. It is based on the concept of Intrinsic Dimension ($I_{d}$) of data, which refers to the minimum number of variables required to describe a dataset. The key idea behind $I_{d}$or is that if there is a correlation between two datasets (i.e., the information from one can partially reconstruct the other), the combined $I_{d}$ of the datasets will be lower than the sum of their individual $I_{d}$s. This method provides both a correlation coefficient and a P-value, obtained through a permutation test, indicating the probability of no correlation.
@@
Finally, Cosine Similarity is another method for measuring similarity between high-dimensional vectors, defined as the cosine of the angle between two vectors. Unlike other methods, cosine similarity is not strictly a measure of correlation. It offers a localised measure of similarity for individual pairs of vectors rather than evaluating entire datasets (i.e., multiple samples of a random variable). Additionally, cosine similarity requires that both vectors exist in the same vector space, making it unsuitable for vectors with different dimensionalities.
IV Experimental results
IV-A Data